OpenStudy (anonymous):

Find the y-intercept of the line whose equation is 2x - 3y = -6.

OpenStudy (anonymous):

would it be 2

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

An easy way to find y intercept is to put the equation into slope intercept form. y=mx+b where b is the y-intercept. Subtract 2x from both sides to give -3y=-2x-6 Divide by -3 to give y=2/3x+2. So yes, the answer is 2 for the y-intercept.